# Multi-phase-field Models of Biological Tissues

### v1

@@ Abstract
The development of complex multicellular organisms from a single parent cell is a highly orchestrated process that cells conduct collectively without central guidance, creating intricate dynamic patterns essential for development and regeneration. Despite significant advances in imaging spatiotemporal dynamics of cell collectives and mechanical characterization techniques, the role of physical forces in biological functions remains poorly understood. Physics-based models are crucial in complementing experiments, providing high-resolution spatiotemporal fields in three dimensions. This review focuses on dense, soft multicellular systems, such as tissues, where mechanical deformation of one cell necessitates the re-organization of neighboring cells. The multi-phase-field model offers a rich physics-based framework to advance our understanding of biological systems and provides a robust playground for non-equilibrium physics of active matter. We discuss the foundational aspects of the multi-phase-field model and their applications in understanding physics of active matter. We also explore the integration of biological physics with experimental data, covering cell migration, heterogeneous cell populations, and confined systems. Finally, we highlight current trends, the importance of multi-phase-field models in biological and physics research, and future challenges.
@@ 1 Introduction
This review is organized as follows: We begin with an introduction to physical models for cell collectives, highlighting the importance of understanding collective cell behavior in development and regeneration. Next, we delve into the multi-phase-field model, discussing its foundational aspects, including time-scale separability, free energy functionals for passive interactions, and non-equilibrium behavior driven by active processes. We then explore the physics of active matter, focusing on collective self-organization, topological defects, active stress chains, and the emergence of biological phenomena. In the subsequent section, we bridge biological physics with experimental integration, covering quantitative modeling informed by experiments, cell migration in 2D and 3D, heterogeneous cell populations, and confined systems. Finally, we conclude with a discussion of current trends, the importance of multi-phase-field models in biological and physics research, and future challenges. Throughout, we emphasize the role of physics-based models in advancing our understanding of multicellular systems, while noting that this review does not address single-cell modeling, molecular-level dynamics, or non-biological applications of multi-phase-field models.
@@ 1.1 Physical models for cell collectives
Although this review focuses on the latest advances in multi-phase-field models and their application to living cells, it is worth briefly discussing other types of approaches. A more comprehensive overview can be found elsewhere [16, 17, 18]: (a) Active network models include vertex-based [19] and Voronoi-based [20] approaches that, in most cases, represent a confluent layer through geometric description of each cell and the associated effective energy functional for the work required to deform that cell. Two-dimensional active network models have played a key role in explaining some of the fundamental biological observations [21, 22] with recent advances extending the model into the third dimension [23] and more complex force generation mechanisms [24]. (b) Particle-based models for living cells are based on granular physics concepts. These models have been instrumental in studying collective motion emerging from local interactions and force generation mechanisms [25, 26, 27]. However, this approach cannot capture cell shape changes (deformation), which is critical for modeling dense, deformable multicellular assemblies. (c) Cellular Potts models [28] represent another paradigm for modeling cellular systems where each cell is defined as a spin collection. These models can provide a detailed description of cell shapes. However, these shapes can be affected by artifacts due to the lattice choice. Furthermore, shape fluctuations depend on a temperature parameter that cannot be directly calibrated against experimental measurements. (d) Continuum models [29, 30, 31, 32, 33, 34, 35] typically leverage hydrodynamics and have been instrumental in modeling the behavior of cell collectives. However, they provide a coarse-grained description of an inherently discrete system and are thus unable to resolve individual cell shapes as well as local force, and density fluctuations that are now understood to be fundamental to many biological functions. Such insights are less likely to be obtained from mean-field descriptions of relevant fields.
The multi-phase-field model overcomes some of these limitations. It can naturally handle deformable interfaces and efficiently scale to a large system in three dimensions, a necessity for understanding inherently three-dimensional biological processes such as embryogenesis and tumorigenesis. Furthermore, its discrete nature, at the scale of an individual cell, can capture large fluctuations while providing a robust way to model proliferation, genetic mutations and/or expressions. This framework is also rather flexible in incorporating various passive and active interactions in a concise and consistent manner. As such, the multi-phase-field model provides a robust framework to establish an integrated understanding of living matter as an interplay between mechanical cues, genetics, and biochemical signals.
@@ 2 Multi-phase-field model
Early phase-field models for modeling cells focused on single and multi-component vesicles [36, 37] and single cell morpho-dynamics [38, 39]. This was followed by models for cell collectives [40, 41, 42, 43] including a number of three-dimensional multi-phase-field models that offer unprecedented insights into the physics of cell collectives [44, 45, 46, 47].
@@ 2.3 Non-equilibrium behavior: active drive
Such local injection of energy breaks time-reversal symmetry and detailed balance, giving rise to fascinating collective phenomena at length scales much larger than a single cell [65], including intricate and dynamic patterns critical for embryogenesis. To this end, multicellular assemblies are a textbook example of active matter, defined as natural or synthetic systems composed of entities that dissipate energy to perform mechanical work on themselves and their environment [66, 32, 67]. As such, living cells can be viewed as a particular instance of active matter with the ability to proliferate, differentiate, and mutate.
@@
Over the past two decades, extensive research has established a strong analogy between living matter – at multiple length scales – and active liquid crystals [70, 71, 72]. This provides a rigorous theoretical framework to understand living matter and, in particular, dynamics of self-organization and collective behavior of cells. Liquid crystals are an intermediate phase of matter between solids and liquids, characterized by a short-range translational order and a quasi-long-range orientational order [73, 74]. There are different phases of liquid crystals; generally named $p$-atic liquid crystals exhibiting $p$-fold rotational symmetry, i.e. symmetry with respect to rotations by $2\pi/p$. This forms the basis for the physics-based description of active force generation at the scale of a single cell and its interplay with the emergence of order at a much larger scale.
@@
where $\zeta_{\text{Q}}$ represents the strength, contractile for $\zeta_{\text{Q}}<0$ and extensile for $\zeta_{\text{Q}}>0$, and $\bm{Q}_{i}=2\left(\vec{n}_{i}\otimes\vec{n}_{i}-\frac{\mathbbm{I}}{2}\vec{n}_{i}^{2}\right)$, where $\mathbbm{I}$ is rank 2 identity tensor. This active nematic stress introduces a dipolar force density along cells’ interfaces (Fig. 1d). The angle associated with $\vec{n}_{i}$ can follow its own dynamics [79] or it can be set equal to cell polarity, $\vec{n}_{i}=\vec{p}_{i}$, e.g. ensuring contractile stresses act in the same direction as protrusion formation. The nematic order along the direction $\vec{n}$ is head-tail symmetric, i.e. $\vec{n}\rightarrow-\vec{n}$. This implies that polar entities with broken front-back symmetry can self-organize into nematic ordering if on average they align together while not pointing to a particular direction.
Another rank 2 tensor construct for active drive focuses on intercellular force generation through adherens junctions [41]. This approach utilizes a shape tensor, $\bm{S}_{i}$, to characterize shape anisotropy such that the direction associated with its largest eigenvalue captures the axis of elongation:
@@
The interplay of emerging order and the nature of active force generation remains an intense area of research [34, 80, 22, 35, 81].
Lastly, using the definitions for active and passive forces (Eqs. (9) and (3), respectively) a rank 2 stress tensorial field can be constructed, providing access to in-plane and out-of-plane stress components. This is best achieved using a discrete stress definition consistent with those used in the granular physics and molecular dynamics communities [82, 83]. Such a definition respects the discrete nature of a multi-phase-field model. For passive stresses another approach can be utilized based on Korteweg stresses for diffusive interfaces [84], commonly used in multiphase fluid mechanics for capillary stresses [85, 86], most recently introduced comprehensively in the contect of a multi-phase-field model [87].
@@
3 Multi-phase-field models and physics of active matter
@@ 3.1 Collective self-organization and emergence of order
The order that emerges from self-organization in living cells governs the most fundamental biological processes, including development. Focusing on multicellular systems as opposed to subcellular ones, nematic order ($p=2$) and the associated half-integer topological defects have been reported in a number of cellular systems [92], including but not limited to Madine-Darby Canine Kidney (MDCK) cells [93, 94], human breast cancer cells [95] and neural progenitor stem cells [96]. Similarly, multi-phase-field models have shown the emergence of nematic order due to force generation mechanisms defined by rank 2 active stress tensors [41, 79] (Eqs. (8) and (7)). Concurrently, the co-emergence of nematic (Fig. 2b) and hexatic orders from active polar forces, Eqs. (5) and (6), has been captured by a multi-phase-field model [46]. Furthermore, Armengol-Collado et al. have demonstrated co-existence of both nematic and hexatic orders in confluent cell layers depending on the length scales of interest: two-fold nematic ($p=2$) is dominant at larger scales, whereas six-fold hexatic ($p=6$) at smaller length scales [35] (Fig. 2c). This is achieved by complementing in vitro experiments on confluent MDCK cells with multi-phase-field and self-propelled voronoi models. This hierarchical structure has also been recovered from a hydrodynamics description of confluent epithelial monolayers [97], hinting at the generality of this approach. Recent experiments complemented with vertex-based modeling show how correlated cell division can lead to the emergence of tetratic ($p=4$) order [22]. Viewing multicellular assemblies as multiscale $p$-atic active liquid crystals provides a powerful framework for understanding the dynamics of self-organization and collective behavior of cells with important implications for $p-$atic defect dynamics that is crucial to some of the most fundamental biological processes.
@@ 3.2 Collective self-organization and breakdown of order: topological defects
Topological defects refer to singularities that disrupt field symmetry. In living cells, they can be formed due to geometric frustration, boundary conditions and/or activity responsible for driving the system away from thermodynamic equilibrium. One of the key features of active liquid crystalline systems is the continuous formation and annihilation of topological defects. The multiphase-field models with shape-determined activity [41] and nematic dipolar activity [79] not only have reproduced the dynamics of topological defects but also accurately captured the flow fields and stresses around the defects (Fig. 2d). This holds for both extensile and contractile systems. Furthermore, Zhang et al. incorporated polar forces that arise from cytoskeletal propulsion, observing a sharp transition from jammed to liquid states, as well as flocking [98].
@@ 3.3 Active stress chains and force transmission
The spatiotemporal dynamics of mechanical stresses play an important but poorly understood role in gene expression [105], genomic damage and potential mutation [106], dynamics of expanding populations [107] as well as cell differentiation [108]. At the same time, the transition of mechanical information is a primary suspect for facilitating collective self-organization, on a length scale much larger than an individual cell, crucial for development and regeneration [109, 12]. In this vein, application of a three-dimensional multi-phase-field model along with extensive in vitro experiments have uncovered how local stress fluctuations modulated via adherens junctions govern some of the most fundamental biological processes such as the fate of an extruding cell [56] and the outcome of mechanical cell competition [110]. These studies highlight the important role intercellular force transmission plays in most vital biological processes such as morphogenesis and cancer invasion. Another recent computational study investigated solid-like to liquid-like transition in active monolayers by focusing on active stress chains and emerging stress patterns [90] (Fig. 2g). By mapping this transition onto the 2D random percolation universality class using two independent active derives, this work revealed the short-range nature of stress correlation near this transition, providing a new context for understanding the non-equilibrium physics of active systems and its connections to glass, jamming and two-dimensional melting transitions. Recent research has shown that activity-induced annealing can lead to a ductile-to-brittle transition in amorphous solids, which parallels the behavior observed in biological tissues under mechanical stress [111]. This insight is crucial for understanding how mechanical forces influence tissue integrity and failure.
3.4 Emerging biological phenomena
@@ 4 Biological physics of cells: integration of experiments with multi-phase-field modeling
Due to the ability to incorporate changes in shape, cell-cell, and cell-substrate interactions, as well as subcellular details, phase-field models have provided valuable theoretical predictions of the biophysics of single cells that have been tested experimentally (reviewed in [16]). In this section, we will focus on the role of multi-phase-field models on the interdisciplinary advances in multicellular systems, starting from parametrization of 2D cell monolayers to 3D models of embryogenesis. We focus particularly on problems where multi-phase-field modeling has been combined and studied together with relevant biological experiments. To bridge these different scales, we first discuss the importance of quantitative parametrization before examining the emergent behaviors that arise from mechanical interactions.
@@ 4.3 Experimentally-informed modeling of collective cell migration in 2D
4.4 From 2D to 3D: cell extrusion from monolayers
@@ 4.7 Confined cellular systems in 3D
5 Conclusion
5.1 Summary of current trends and contributions
Multi-phase-field models have become a pivotal tool in understanding the complex behaviors of cells and tissues and the emergent dynamics of deformable active particles. These models have evolved to incorporate a wide range of physical principles, enabling the simulation of intricate biological processes. Recent advancements have focused on enhancing the accuracy and applicability of these models in biological contexts. For instance, the emergence of active nematic behavior in cell monolayers has been demonstrated, highlighting the role of cell polarity and mechanical interactions in tissue dynamics [41, 102]. Notably, integrating substrate interactions and mechanical constraints has uncovered how cellular activity generates collective flow patterns, shedding light on how tissues self-organize through interacting with their underlying environments [46].
Incorporating cell-cell adhesion and cortical tension in phase-field models has led to a better understanding of collective cell migration and tissue surface tension [136]. The influence of microscopic details on collective cell behavior has been explored, showing how variations in cell shape and intercellular forces can lead to different emergent phenomena [79, 43]. Moreover, these models have facilitated the study of contact inhibition and jamming transitions, revealing how mechanical constraints regulate tissue fluidity and cell proliferation. Mechanochemical interactions underlying cell motility have been modeled to provide a comprehensive framework for understanding cell migration in complex environments [137]. Integration of stochastic hydrodynamics and reaction-diffusion modeling has captured the dynamic processes of cell signaling and chemotaxis [17].
Recent studies have further enriched our understanding of these models. For instance, intercellular friction and motility drive orientational order in cell monolayers, elucidating the solid-liquid transition and the emergence of nematic and hexatic orders [102]. The role of force transmission in mechanical cell competition has been highlighted, revealing how differences in intercellular adhesion can determine the outcome of cell competition [110]. Additionally, these models have revealed how local cellular compliance and force fluctuations near tissue interfaces influence the mechanical resilience of cell collectives, offering insights into morphogenetic robustness. Local yield and compliance in active cell monolayers have been investigated, demonstrating how cell motility and deformability influence the rheological properties of tissues [138].
Furthermore, reaction-diffusion processes and energy minimization have been proposed as mechanisms to sense cell shape, providing a deeper understanding of cell polarity and patterning [139]. The study of stress chains near the solid-to-liquid transition in active monolayers has revealed short-range stress correlations, emphasizing the role of localized stress accumulation in mediating tissue remodeling and phase transitions in active systems [90].
By integrating these diverse contributions, multi-phase-field models have evolved into a powerful framework for predicting and analyzing emergent behaviors in biological systems. As interdisciplinary collaborations grow, these models will likely serve as a bridge between physics-based modeling and experimental tissue engineering, enabling quantitative predictions of tissue development and disease progression.
5.2 Importance of multi-phase-field models in biological research
The significance of multi-phase-field models in biological research cannot be overstated. These models offer a versatile and robust framework for simulating the intricate behaviors of cells and tissues, bridging the gap between theoretical predictions and experimental observations. By incorporating key physical principles such as cell adhesion, motility, and mechanical interactions, multi-phase-field models enable researchers to explore the fundamental principles governing tissue dynamics and development.
The ability to simulate both individual cell behaviors and collective phenomena makes these models invaluable for studying a wide range of biological systems, from epithelial layers to tumor growth. The insights gained from these simulations can inform experimental designs, guide the development of new therapeutic strategies, and enhance our understanding of disease progression. As computational power continues to grow, the integration of machine learning with multi-phase-field models could further refine parameter inference, enabling real-time simulations of complex cellular behaviors.
Recent advancements have underscored the importance of these models. For example, the study of dynamic forces shaping the survival fate of eliminated cells has revealed how modified force transmission across adherens junctions can inhibit apoptotic cell eliminations, linking intercellular force transmission to cell extrusion mechanisms [56].
Moreover, the critical role of force transmission in mechanical cell competition has been revealed, showing how differences in intercellular adhesion can determine the outcome of cell competition [110]. This has important implications for understanding tissue homeostasis and the development of diseases such as cancer. Additionally, phase-field models have provided a framework for investigating mechanical feedback loops that influence stem cell differentiation, revealing how microenvironmental cues dictate lineage specification.
The importance of confinement and cytokinesis machinery in promoting cell dissociation from collectively migrating strands has been highlighted, emphasizing the role of mechanical forces in cell migration and having important implications for understanding processes such as wound healing and cancer metastasis [113].
In long term, with growing computational power and as biological research moves towards personalized medicine, the predictive power of models such as multi-phase-field framework in reconstructing specific tissue morphologies could play a role in developing tailored therapeutic strategies.
5.3 Importance of multi-phase-field models in physics research
One of the key areas where multi-phase-field models have made significant contributions is in understanding the collective behaviors of cells and tissues. These models have been used to study how mechanical interactions and active forces drive the organization and dynamics of cell collectives. For example, recent research has shown that activity-induced annealing can lead to a ductile-to-brittle transition in amorphous solids, which parallels the behavior observed in biological tissues under mechanical stress [111]. This analogy between biological and non-living matter provides an opportunity to apply concepts from condensed matter physics to tissue mechanics, offering new perspectives on the failure modes of biological structures.
multi-phase-field models have also been instrumental in studying the emergent properties of active matter. Active matter systems, such as bacterial colonies, cell monolayers, and cytoskeletal networks, exhibit complex behaviors that arise from the interactions between individual active units. These models have provided insights into phenomena such as flocking, swarming, and collective migration.
The ability to simulate the evolution of interfaces and the interactions between different phases makes multi-phase-field models invaluable for understanding the physics of living systems. For instance, recent studies have demonstrated that stress accumulation at cellular interfaces can generate force chains, akin to granular matter, revealing how active cell collectives maintain structural integrity under compression.
Furthermore, multi-phase-field models have been applied to study cell-cell sorting and competition, which are fundamental processes in tissue development and homeostasis. By simulating the interactions between different cell types, researchers can gain insights into how differences in cell adhesion, motility, and mechanical properties drive cell sorting and competition [140].
Capturing 3D shape changes of cells in collectives is another critical area where multi-phase-field models have shown promise. These models can simulate the complex shape changes that cells undergo during processes such as migration, division, and differentiation, providing valuable insights into the mechanics of tissue morphogenesis [141].
The growing use of phase-field models in connection with diverse experimental setups suggests that these frameworks could serve as computational testbeds for designing biomaterials with tunable mechanical properties, offering a bridge between theoretical physics and biomedical applications.
Recent studies have highlighted the role of cell deformability in driving fluid-to-fluid phase transitions in active cell monolayers, providing new insights into the rheological properties of tissues and their response to mechanical forces [142]. The study of T1 transitions, or cell intercalations, has revealed their importance in tissue fluidization and mixing, further emphasizing the relevance of multi-phase-field models in understanding the physics of living matter [143]. These transitions, characterized by neighbor exchanges, generate transient saddle-point flows and localized bursts in cell speed, which are crucial for tissue remodeling and development. Understanding these processes at a quantitative level is essential not just for deciphering fundamental biological mechanisms but also for designing bioengineered tissues and improving medical interventions. By leveraging multi-phase-field models, it is possible to predict how mechanical forces shape tissues, opening pathways to controlling morphogenesis, optimizing regenerative medicine approaches, and even tackling disease progression in mechanically driven disorders.
5.4 Challenges and future directions

### v2

@@ Abstract
Understanding how cells coordinate their behaviors to produce large-scale patterns and functions is central to deciphering biological processes ranging from tissue development and regeneration to cancer progression and morphogenesis. Despite advances in imaging and mechanical characterization, the role of physical forces in collective cell dynamics remains incompletely understood. Physics-based models are essential for complementing experimental data, offering access to high-resolution spatiotemporal fields and enabling mechanistic insights into complex multicellular systems. This review focuses on dense, soft tissues, where the mechanical deformation of one cell drives re-organization of its neighbors, giving rise to emergent behaviors such as orientational order and long-range force transmission. The multi-phase-field model provides a powerful and versatile framework to investigate such systems, bridging biological phenomena and the non-equilibrium physics of active matter. We discuss the theoretical foundations of the model and its applications to a range of biological contexts, including cell migration, heterogeneous populations, confined geometries, and metastasis. We also emphasize the integration of simulations with experimental data, highlighting how this approach is reshaping our understanding of tissue mechanics, collective order, and force transmission. Finally, we outline current trends and future challenges in applying multi-phase-field models to biology and soft matter physics.
@@ 1 Introduction
This review is organized as follows: We begin with an introduction to physical models for cell collectives, highlighting the importance of understanding collective cell behavior in development and regeneration. Next, we delve into the multi-phase-field model, discussing its foundational aspects, including time-scale separability, free energy functionals for passive interactions, and non-equilibrium behavior driven by active processes. We then explore the physics of active matter, focusing on collective self-organization, topological defects, active stress chains, and the emergence of biological phenomena. In the subsequent section, we bridge biological physics with experimental integration, covering quantitative modeling informed by experiments, cell migration in 2D and 3D, heterogeneous cell populations, and confined systems. Finally, we conclude with a discussion of current trends, the importance of multi-phase-field models in biological and physics research, and future challenges.
Due to the ability to incorporate changes in shape, cell-cell, and cell-substrate interactions, as well as subcellular details, phase-field models have provided valuable theoretical predictions of the biophysics of single cells that have been tested experimentally (reviewed in [7]). Throughout, we emphasize the role of physics-based models in advancing our understanding of multicellular systems, while noting that this review does not address single-cell modeling, molecular-level dynamics, or non-biological applications of multi-phase-field models.
@@ 1.1 Physical models for cell collectives
Although this review focuses on the latest advances in multi-phase-field models and their application to living cells, it is worth briefly discussing other types of approaches. A more comprehensive overview can be found elsewhere [7, 17, 18, 19]: (a) Active network models include vertex-based [20] and Voronoi-based [21] approaches that, in most cases, represent a confluent layer through geometric description of each cell and the associated effective energy functional for the work required to deform that cell. Two-dimensional active network models have played a key role in explaining some of the fundamental biological observations [22, 23] with recent advances extending the model into the third dimension [24] and more complex force generation mechanisms [25]. (b) Particle-based models for living cells are based on granular physics concepts. These models have been instrumental in studying collective motion emerging from local interactions and force generation mechanisms [26, 27, 28]. However, this approach cannot capture cell shape changes (deformation), which is critical for modeling dense, deformable multicellular assemblies. (c) Cellular Potts models [29] represent another paradigm for modeling cellular systems where each cell is defined as a spin collection. These models can provide a detailed description of cell shapes. However, these shapes can be affected by artifacts due to the lattice choice. Furthermore, shape fluctuations depend on a temperature parameter that cannot be directly calibrated against experimental measurements. (d) Continuum models [30, 31, 32, 33, 34, 35, 36] typically leverage hydrodynamics and have been instrumental in modeling the behavior of cell collectives. However, they provide a coarse-grained description of an inherently discrete system and are thus unable to resolve individual cell shapes as well as local force, and density fluctuations that are now understood to be fundamental to many biological functions. Such insights are less likely to be obtained from mean-field descriptions of relevant fields.
The multi-phase-field model overcomes some of these limitations It can naturally handle deformable interfaces and efficiently scale to a large system in three dimensions, a necessity for understanding inherently three-dimensional biological processes such as embryogenesis and tumorigenesis. Furthermore, its discrete nature, at the scale of an individual cell, can capture large fluctuations while providing a robust way to model proliferation, genetic mutations and/or expressions. This framework is also able to represent an arbitrary level of confluence [37, 38, 39] and rather flexible in incorporating various passive and active interactions in a concise and consistent manner. As such, the multi-phase-field model provides a robust framework to establish an integrated understanding of living matter as an interplay between mechanical cues, genetics, and biochemical signals.
@@ 2 Multi-phase-field model
Early phase-field models for modeling cells focused on single and multi-component vesicles [40, 41] and single cell morpho-dynamics [42, 43, 44, 45]. This was followed by models for cell collectives [37, 46, 47, 48, 49] including a number of three-dimensional multi-phase-field models that offer unprecedented insights into the physics of cell collectives [50, 51, 52, 53].
@@ 2.3 Non-equilibrium behavior: active drive
Such local injection of energy breaks time-reversal symmetry and detailed balance, giving rise to fascinating collective phenomena at length scales much larger than a single cell [71], including intricate and dynamic patterns critical for e.g. embryogenesis. To this end, multicellular assemblies are a textbook example of active matter, defined as natural or synthetic systems composed of entities that dissipate energy to perform mechanical work on themselves and their environment [72, 33, 73]. As such, living cells can be viewed as a particular instance of active matter with the ability to proliferate, differentiate, and mutate.
@@
where $\zeta_{\text{Q}}$ represents the strength, contractile for $\zeta_{\text{Q}}<0$ and extensile for $\zeta_{\text{Q}}>0$, and $\bm{Q}_{i}=2\left(\vec{n}_{i}\otimes\vec{n}_{i}-\frac{\bm{I}}{2}\vec{n}_{i}^{2}\right)$, where $\bm{I}$ is rank 2 identity tensor. This active nematic stress introduces a dipolar force density along cells’ interfaces (Fig. 1d). The angle associated with $\vec{n}_{i}$ can follow its own dynamics [39] or it can be set equal to cell polarity, $\vec{n}_{i}=\vec{p}_{i}$, e.g. ensuring contractile stresses act in the same direction as protrusion formation. The nematic order along the direction $\vec{n}$ is head-tail symmetric, i.e. $\vec{n}\rightarrow-\vec{n}$. This implies that polar entities with broken front-back symmetry can self-organize into nematic ordering if on average they align together while not pointing to a particular direction.
Another rank 2 tensor construct for active drive focuses on the generation of intercellular force through the adherent junctions [47]. This approach utilizes a shape tensor, $\bm{S}_{i}$, to characterize shape anisotropy such that the direction associated with its largest eigenvalue captures the axis of elongation:
@@
The interplay of emerging order and the nature of active force generation remains an intense area of research [35, 80, 23, 36, 81], which we will discuss further in Section 3.
Lastly, using the definitions for active and passive forces (Eqs. (9) and (3), respectively) a rank 2 stress tensorial field can be constructed, providing access to in-plane and out-of-plane stress components. This is best achieved using a discrete stress definition consistent with those used in the granular physics and molecular dynamics communities [82, 83]. Such a definition respects the discrete nature of a multi-phase-field model. For passive stresses another approach can be utilized based on Korteweg stresses for diffusive interfaces [84], commonly used in multiphase fluid mechanics for capillary stresses [85, 86], most recently introduced in the context of a multi-phase-field model [87].
3 Multi-phase-field models and physics of active matter
@@ 3.1 Collective self-organization and emergence of order
Over the past two decades, extensive research has established a strong analogy between living matter – at multiple length scales – and active liquid crystals [92, 93, 94]. This provides a rigorous theoretical framework to understand living matter and, in particular, dynamics of self-organization and collective behavior of cells and how it affects biological processes such as cancer metastasis [95]. Liquid crystals are an intermediate phase of matter between solids and liquids, characterized by a short-range translational order and a quasi-long-range orientational order [96, 97]. There are different phases of liquid crystals; generally named $p$-atic liquid crystals exhibiting $p$-fold rotational symmetry, i.e. symmetry with respect to rotations by $2\pi/p$. This forms the basis for the physics-based description of active force generation at the scale of a single cell and its interplay with the emergence of order at a much larger scale.
The order that emerges from self-organization in living cells governs the most fundamental biological processes such as development. Focusing on multicellular systems as opposed to subcellular ones, nematic order ($p=2$) and the associated half-integer topological defects have been reported in a number of cellular systems [98], including but not limited to Madine-Darby Canine Kidney (MDCK) cells [99, 100], human breast cancer cells [101] and neural progenitor stem cells [102]. Similarly, multi-phase-field models have shown the emergence of nematic order due to force generation mechanisms defined by rank 2 active stress tensors [47, 39] (Eqs. (8) and (7)). Concurrently, the co-emergence of nematic (Fig. 2b) and hexatic orders from active polar forces, Eqs. (5) and (6), has been captured by a multi-phase-field model [52]. Furthermore, Armengol-Collado et al. have demonstrated co-existence of both nematic and hexatic orders in confluent cell layers depending on the length scales of interest: two-fold nematic ($p=2$) is dominant at larger scales, whereas six-fold hexatic ($p=6$) at smaller length scales [36] (Fig. 2c). This is achieved by complementing in vitro experiments on confluent MDCK cells with multi-phase-field and self-propelled voronoi models. This hierarchical structure has also been recovered from a hydrodynamics description of confluent epithelial monolayers [103], hinting at the generality of this approach. Recent experiments complemented with vertex-based modeling show how correlated cell division can lead to the emergence of tetratic ($p=4$) order [23]. Viewing multicellular assemblies as multiscale $p$-atic active liquid crystals provides a powerful framework for understanding the dynamics of self-organization and collective behavior of cells with important implications for $p-$atic defect dynamics that is crucial to some of the most fundamental biological processes.
@@ 3.2 Collective self-organization and breakdown of order: topological defects
Topological defects refer to singularities that disrupt field symmetry. In living cells, they can be formed due to geometric frustration, boundary conditions and/or activity responsible for driving the system away from thermodynamic equilibrium. One of the key features of active liquid crystalline systems is the continuous formation and annihilation of topological defects. The multi-phase-field models with shape-determined activity [47] and nematic dipolar activity [39] not only have reproduced the dynamics of topological defects but also accurately captured the flow fields and stresses around the defects (Fig. 2d). This holds for both extensile and contractile systems. Furthermore, Zhang et al. incorporated polar forces that arise from cytoskeletal propulsion, observing a sharp transition from jammed to liquid states, as well as flocking [104].
@@ 3.3 Active stress chains and force transmission
The spatiotemporal dynamics of mechanical stresses play an important but poorly understood role in gene expression [111], genomic damage and potential mutation [112], dynamics of expanding populations [113] as well as cell differentiation [114]. At the same time, the transition of mechanical information is a primary suspect for facilitating collective self-organization, on a length scale much larger than an individual cell, crucial for development and regeneration [115, 13]. A recent computational study investigated solid-like to liquid-like transition in active monolayers by focusing on active stress chains and emerging stress patterns [90] (Fig. 2g). By mapping this transition onto the 2D random percolation universality class using two independent active derives, this work revealed the short-range nature of stress correlation near this transition, providing a new context for understanding the non-equilibrium physics of active systems and its connections to glass, jamming and two-dimensional melting transitions. Recent research has shown that activity-induced annealing can lead to a ductile-to-brittle transition in amorphous solids, which parallels the behavior observed in biological tissues under mechanical stress [116]. This insight is crucial for understanding how mechanical forces influence tissue integrity and failure. In Section 4, focused on biological physics, we will thoroughly discuss the role of force transmission on biological processes such as the fate of an extruding cell and cell competition.
@@ 4 Biological physics of cells: integration of experiments with multi-phase-field modeling
In this section, we will focus on the role of multi-phase-field models on the interdisciplinary advances in multicellular systems, starting from parametrization of 2D cell monolayers to 3D models of embryogenesis. We focus particularly on problems where multi-phase-field modeling has been combined and studied together with relevant biological experiments. To bridge these different scales, we first discuss the importance of quantitative parametrization before examining the emergent behaviors that arise from mechanical interactions.
@@ 4.3 Experimentally-informed modeling of collective cell migration in 2D
4.4 From 2D to 3D: mechanical imprints of live versus dead cell elimination
@@ 4.7 Confined cellular systems in 3D
5 Challenges and future directions